3/ Find the inverse of the function f(x) and state the domain and the range of f1. f(x) = 3 cos (5x) + 2 for 0 sxsn/5
Answer #1

y = f(x) = 360$ (5w) #2 for 0<x<T For x in the given range, ye for decreases morbonically with increase in t because cosine fy = f(x) = 360$ (5w) #2 for 0<x<T For x in the given range, ye for decreases morbonically with increase in t because cosine fFor X:-1, f(x) = f Cos (-1-2) . : - Cos? (-1) = . We restrict the value of arccos function to Qi (5 m fer ve arguments & Co